OpenStudy (anonymous):

Let f be a function such that f(-2)=5, f(0)=-2, and f(1)=0. Give the coordinates of three points on the graph of y=f(x)+1

jimthompson5910 (jim_thompson5910):

y=f(x)+1 y=f(-2)+1 y=5+1 y = 6 So (-2,6) is one point on y = f(x)+1

OpenStudy (anonymous):

Thanks!